# Whose Facts Win? LLM Source Preferences under Knowledge Conflicts

### v1

@@ 3.2 Synthetic Sources
We collect all U.S. newspaper names from Media Bias/Fact Check [^3] https://mediabiasfactcheck.com/, mask all location names using SpaCy (Honnibal et al., 2020), and extract the $150$ most frequent 22-, 33- and 44-grams. After deduplication, $59$ newspaper templates such as "The {LOC} Herald" remain. We fill these templates with fictional locations from NeoQA to create synthetic newspaper names.
@@ 3.3 Evaluation Method
A positive value indicates that $x$ increases the support for option AA more than $y$ supports $B$. We aggregate source preferences $\widehat{SP}$ for source types $X,Y$ over a dataset $D$ of conflict pairs by averaging SP over $D$ and drawing instances of $X$ and $Y$ for every pair of conflicting tables:
@@ 4.2 Intra-Type Source Preference Behavior
We have shown that models display clear source preferences when resolving knowledge conflicts. Next, we test whether source preferences can be obtained via context-independent prompting to select the more credible source, and to what extent prompted preferences agree with behavioral preferences. Differences between prompting and behavior have also been studied in stereotypes, grammaticality, and value alignment (Goldfarb-Tarrant et al., 2021; Hu and Levy, 2023; Shen et al., 2025a). Specifically, we elicit credibility judgments on conflicting source pairs with prompts like: "Who would you rate higher in terms of credibility? (A) Melvidra News Press (B) Matthew Chen". For every prior source matchup (except regionality, which requires context), we construct $2,000$ inputs, consisting of $1,000$ source pairs, each paired with a question sampled from 2020 templates, and presented in both orders (see Appendix D for prompts and questions). We report probability deviation from 50% of answer tokens representing each source.
@@ 6 Mitigating Repetition Bias
$+(1-\lambda)$ $D_{KL}\!\left(f_{t}(C^{\prime}_{R})\,\|\,f_{s}(C^{\prime}_{U})\right)$
@@ E.1 Alternative Induction of a Credibility Hierarchy
Instead of direct source match-ups, we can order the four source types by their $\widehat{SP}$ value when in conflict with No source available. For clarity, we will refer to this method as the attribution-based ranking and our main method as the match-up-based ranking. The attribution-based rankings lead to an inter-model Kendall’s $W$ of $0.66$, with 99 out of $13$ models ranking both institutional sources over both personal ones. Comparing attribution-based and matchup-based rankings shows very stable results: For $10$ out of $13$ models, both yield the same ranking, with only minimal differences between them and an average Kendall’s $\tau$ of $0.87$. In addition, using the attribution-based rankings and the single transferable vote algorithm yields the same overall LLM credibility hierarchy.
@@ E.2 Different Answer Options
When comparing these results to our main results of Section 4.1, we see that $12$ out of $13$ models produce the same match-up-based rankings with an average Kendall’s $\tau$ of $0.97$, and 88 out of $13$ models produce the same attribution-based rankings with an average Kendall’s $\tau$ of $0.87$. The source credibility hierarchy induced with the single transferable vote algorithm is the same for both methods and identical to the one from Section 4.1.
@@ E.3 Different Instruction
Once again, when varying the instruction, pairwise match-ups between source types are strictly transitive for all models. Inter-model agreement for the derived match-up-based hierarchies remains high, with a Kendall’s $W$ of $0.71$, and 99 out of $13$ models placing both institutional sources over both personal ones. Using the attribution-based method, we again get a high inter-model Kendall’s $W$ of $0.71$ and 99 of $13$ models preferring institutional sources. 99 out of $13$ models produce the same ranking with both methods. The average Kendall’s $\tau$ when comparing the hierarchies created by the two methods is $0.90$.
When comparing these results to our main results of Section 4.1, we find that $10$ out of $13$ models produce the same match-up-based ranking with an average Kendall’s $\tau$ of $0.90$ and 88 out of $13$ models produce the same attribution-based rankings with an average Kendall’s $\tau$ of $0.87$. The source credibility hierarchy induced with the single transferable vote algorithm is identical for both methods and identical to the one from Section 4.1.
@@ E.4 Prompt with Lower Source Focus
Source preferences in this scenario are shown in Figure 21 and 22. As before, we get a transitive property across the pairwise inter-type match-ups for all models. Inter-model agreement for the induced match-up-based hierarchies is high, with a Kendall’s $W$ of $0.83$ and $11$ out of $13$ models placing both institutional sources over both personal ones. Using the attribution-based method, analogous to our findings with prompt-based mitigation (Section 5), removing every mention of source information in the instruction and from unattributed tables lessens the absolute effect of source preference across models. We still get a high inter-model Kendall’s $W$ of $0.66$ and 99 of $13$ models preferring institutional sources. $10$ out of $13$ models produce the same ranking with both methods. The average Kendall’s $\tau$ when comparing the hierarchies created by the two methods is $0.87$.
Comparing these results to our main results of Section 4.1, we find that 88 out of $13$ models produce the same match-up-based rankings with an average Kendall’s $\tau$ of $0.87$ and only 44 out of $13$ models produce the same attribution-based rankings with an average Kendall’s $\tau$ of $0.72$. Once again, the source credibility hierarchy induced by the single transferable vote algorithm is identical across methods and matches the one in Section 4.1.
@@ Appendix I Credibility Prompting for All Models
We use one Quadro RTX 6000 Nvidia GPU to fine-tune the Gemma-3-4B model. We use a batch size of 88, a learning rate of $2e^{-4}$ with a warm-up period in the first $10\%$ of training steps. We insert and fine-tune $32$ million LoRA parameters with typical hyperparameters of $r=16$, $\alpha=16$ and dropout of $0.05$. The $\lambda$ for weighting loss terms was not extensively optimized, but selected from a search space of $0.5,0.75$ and $0.9$. Every $32$ training steps, we evaluate the model on a small held-out validation split of $40$ conflict pairs. Based on this validation, we employ early stopping with a patience of 22 to avoid overfitting for a maximum of 44 epochs, where training terminates during the second epoch. This is followed up by a $300$ step training epoch, only using the first KL-divergence loss term ($\lambda=1$). The total training time for the used hardware was less than one hour.

### v2

@@ 3.2 Synthetic Sources
We collect all U.S. newspaper names from Media Bias/Fact Check [^3] https://mediabiasfactcheck.com/, mask all location names using SpaCy (Honnibal et al., 2020), and extract the $150$ most frequent $2$-, 33- and $4$-grams. After deduplication, $59$ newspaper templates such as "The {LOC} Herald" remain. We fill these templates with fictional locations from NeoQA to create synthetic newspaper names.
@@ 3.3 Evaluation Method
A positive value indicates that $x$ increases the support for option $A$ more than $y$ supports $B$. We aggregate source preferences $\widehat{SP}$ for source types $X,Y$ over a dataset $D$ of conflict pairs by averaging SP over $D$ and drawing instances of $X$ and $Y$ for every pair of conflicting tables:
@@ 4.2 Intra-Type Source Preference Behavior
We have shown that models display clear source preferences when resolving knowledge conflicts. Next, we test whether source preferences can be obtained via context-independent prompting to select the more credible source, and to what extent prompted preferences agree with behavioral preferences. Differences between prompting and behavior have also been studied in stereotypes, grammaticality, and value alignment (Goldfarb-Tarrant et al., 2021; Hu and Levy, 2023; Shen et al., 2025a). Specifically, we elicit credibility judgments on conflicting source pairs with prompts like: "Who would you rate higher in terms of credibility? (A) Melvidra News Press (B) Matthew Chen". For every prior source matchup (except regionality, which requires context), we construct $2,000$ inputs, consisting of $1,000$ source pairs, each paired with a question sampled from $20$ templates, and presented in both orders (see Appendix D for prompts and questions). We report probability deviation from 50% of answer tokens representing each source.
@@ 6 Mitigating Repetition Bias
$+(1-\lambda)$ $D_{KL}\!\left(f_{t}(C^{\prime}_{U})\,\|\,f_{s}(C^{\prime}_{R})\right)$
@@ E.1 Alternative Induction of a Credibility Hierarchy
Instead of direct source match-ups, we can order the four source types by their $\widehat{SP}$ value when in conflict with No source available. For clarity, we will refer to this method as the attribution-based ranking and our main method as the match-up-based ranking. The attribution-based rankings lead to an inter-model Kendall’s $W$ of $0.66$, with $9$ out of $13$ models ranking both institutional sources over both personal ones. Comparing attribution-based and matchup-based rankings shows very stable results: For $10$ out of $13$ models, both yield the same ranking, with only minimal differences between them and an average Kendall’s $\tau$ of $0.87$. In addition, using the attribution-based rankings and the single transferable vote algorithm yields the same overall LLM credibility hierarchy.
@@ E.2 Different Answer Options
When comparing these results to our main results of Section 4.1, we see that $12$ out of $13$ models produce the same match-up-based rankings with an average Kendall’s $\tau$ of $0.97$, and $8$ out of $13$ models produce the same attribution-based rankings with an average Kendall’s $\tau$ of $0.87$. The source credibility hierarchy induced with the single transferable vote algorithm is the same for both methods and identical to the one from Section 4.1.
@@ E.3 Different Instruction
Once again, when varying the instruction, pairwise match-ups between source types are strictly transitive for all models. Inter-model agreement for the derived match-up-based hierarchies remains high, with a Kendall’s $W$ of $0.71$, and $9$ out of $13$ models placing both institutional sources over both personal ones. Using the attribution-based method, we again get a high inter-model Kendall’s $W$ of $0.71$ and $9$ of $13$ models preferring institutional sources. $9$ out of $13$ models produce the same ranking with both methods. The average Kendall’s $\tau$ when comparing the hierarchies created by the two methods is $0.90$.
When comparing these results to our main results of Section 4.1, we find that $10$ out of $13$ models produce the same match-up-based ranking with an average Kendall’s $\tau$ of $0.90$ and $8$ out of $13$ models produce the same attribution-based rankings with an average Kendall’s $\tau$ of $0.87$. The source credibility hierarchy induced with the single transferable vote algorithm is identical for both methods and identical to the one from Section 4.1.
@@ E.4 Prompt with Lower Source Focus
Source preferences in this scenario are shown in Figure 21 and 22. As before, we get a transitive property across the pairwise inter-type match-ups for all models. Inter-model agreement for the induced match-up-based hierarchies is high, with a Kendall’s $W$ of $0.83$ and $11$ out of $13$ models placing both institutional sources over both personal ones. Using the attribution-based method, analogous to our findings with prompt-based mitigation (Section 5), removing every mention of source information in the instruction and from unattributed tables lessens the absolute effect of source preference across models. We still get a high inter-model Kendall’s $W$ of $0.66$ and $9$ of $13$ models preferring institutional sources. $10$ out of $13$ models produce the same ranking with both methods. The average Kendall’s $\tau$ when comparing the hierarchies created by the two methods is $0.87$.
Comparing these results to our main results of Section 4.1, we find that $8$ out of $13$ models produce the same match-up-based rankings with an average Kendall’s $\tau$ of $0.87$ and only $4$ out of $13$ models produce the same attribution-based rankings with an average Kendall’s $\tau$ of $0.72$. Once again, the source credibility hierarchy induced by the single transferable vote algorithm is identical across methods and matches the one in Section 4.1.
@@ Appendix I Credibility Prompting for All Models
We use one Quadro RTX 6000 Nvidia GPU to fine-tune the Gemma-3-4B model. We use a batch size of $8$, a learning rate of $2e^{-4}$ with a warm-up period in the first $10\%$ of training steps. We insert and fine-tune $32$ million LoRA parameters with typical hyperparameters of $r=16$, $\alpha=16$ and dropout of $0.05$. The $\lambda$ for weighting loss terms was not extensively optimized, but selected from a search space of $0.5,0.75$ and $0.9$. Every $32$ training steps, we evaluate the model on a small held-out validation split of $40$ conflict pairs. Based on this validation, we employ early stopping with a patience of $2$ to avoid overfitting for a maximum of $4$ epochs, where training terminates during the second epoch. This is followed up by a $300$ step training epoch, only using the first KL-divergence loss term ($\lambda=1$). The total training time for the used hardware was less than one hour.